# Supervising 3D Talking Head Avatars with Analysis-by-Audio-Synthesis

### v3

@@ 1 Introduction
In summary, our contributions are: (1) A novel mesh-to-speech model that regresses sound from 3D facial animations. (2) A new form of cross-modal audio-based self-supervision via comparison of the representations of the input speech and the animation-produced output speech. (3) THUNDER, a 3D speech-driven animation method based on diffusion, capable of generating a variety of facial expressions while maintaining accurate lip-sync. The code, data annotations and models will be made publicly available .
@@ 2.2 Silent-video-to-speech (SVTS)
STVS [20] was the first method to produce intelligible audio on large-scale in-the-wild datasets such as LRS2 and LRS3 [1] by regressing spectrograms from mouth crops. The spectrograms are converted to the final waveform with a pretrained vocoder[12]. Follow-up methods improve the prediction quality [48] by adding a surrogate ASR loss or predicting features from an ASR model [12] The most recent architectures are based on diffusion, producing audio often indistinguishable from real speech [11, 96]. These, however, are not suitable for a self-supervised reconstruction loop, due to the iterative nature of diffusion models. Hence, thanks to its simple, yet efficient feed-forward design and a relatively straightforward architecture and training, we base our mesh-to-speech model on Choi et al.[12].
@@ 5.3.1 Quantitative evaluation
Does the input space for M2S matter? While all models are improve lip-sync metrics, mouth2speech has the strongest effect on lip-sync metrics, likely thanks to its localized effect on the mouth region (Tab. 2, experiment 1).
@@ 7.2.1 Experiments on TFHP
Consistently to the experiments on THUNDERSET, we find that THUNDER models benefit from the M2S loss, which can be observed in improvement on all lip-sync metrics. The improvement is more dramatic in THUNDER-F models, but manifests itself also in the models with trainable audio encoders (THUNDER-F). Similarly to our experiments on THUNDERSET, the application of M2S comes with a small reduction in upper-face diversity. Finally, THUNDER-T outperforms DiffPoseTalk an all lip-sync metrics. The metrics can be seen in Tab. 5, experiment 1.
@@
Finally, we train THUNDER models with with the DiffPoseTalk contrastive style features as conditions. The results can be found in See Tab. 5, experiment 3. We find that conditioning with the style vector results in better lip-sync performance, lower face diversity, higher beat alignment and lower pose diversity compared to measurements in See Tab. 5, experiment 2. All of these are consistent with expectations, since passing in a style specification reduces the distribution of possible outcomes. Regardless of the above, the application of M2S has the expected effect - improved lip-sync metrics and lower diversity.
@@
A recent paper by Chae et al. [8] proposes a new self-supervised speech-mesh representation, in form of a bi-modal masked autoencoder. The architecture consists of two encoders and two decoders (one for the the mesh modality and one for the audio modality). The model is trained with two loss terms - the autoencoder reconstruction losses for both modalities and the InfoNCE loss which brings the two modalities into one coherent features space. This is similar to what CLIP [66] does with images and text. The authors show that their representation can be used in training other speech-driven animation methods. They use their frozen autoencoder in training, passing the input audio and output geometry in, extracting the features for both. Then, they apply the same InfoNCE loss between the two modalities and use it to co-supervise the talking head avatar system. In this experiment, we incorporate this perceptual loss into the THUNDER training. The training loss is the given as:
@@
where $w_{perc}\mathcal{L}_{perc}$ is the perceptual term. We follow the authors and the weight accordingly to . For completeness, we also report the new metrics proposed by Chae et al. [8] - mean temporal misalignment (MTM) and Perceptual Lip Readability Score (PLRS) along with the other metrics used in this paper. MTM measures the temporal misalignment using a Derivative Dynamic Time Warping (DDTW). SLRS is the same InfoNCE-based perceptual loss used for training, but computed by a different instance of the same architecture. The authors released two instances - one to be used as the perceptual loss and one to be used for evaluation with SLRS. We utilize both of the speech-mesh autoencoder models released by the authors accordingly. Tab. 5, experiment 4 reports the results of this experiment. In this experiment we investigate the following:
@@
(b) Does training our diffusion-based speech-driven animation method be trained with this perceptual loss? Yes, THUNDER models trained without mesh-to-speech but with the perceptual loss improve considerably on PLRS. This is expected since PLRS is computed by a model with the same architecture as the network which provides the perceptual loss function. We also observe improvement on MTM. Furthermore, we observe improvement on our lip-sync related metrics in models with the frozen backbone (-F). Models with trainable backbones (-T) do not seem to improve on LVE, CCC, PCC or DTW when the perceptual loss is applied.
(c) Can THUNDER be trained with both mesh-to-speech and the perceptual loss. Yes. Both losses can be applied together. The mesh-to-speech loss has a strong impact on LVE, CCC, PCC, DTW and MTM, while the perceptual loss strongly affects PLRS, MTM and with -F models also improves CCC and PCC. THUNDER-F with both mesh-to-speech and the perceptual loss appears to get the best of both worlds, while THUNDER-T with mesh-to-speech only is competetive with THUNDER-T with both losses. In any case, applying the mesh-to-speech is beneficial.
@@ 7.2.3 Extended Ablation and Sensitivity Analysis
Table 7: Ablation study and sensitivity analysis. Here we analyze the effect of the mesh-to-speech loss on the talking head avatar training. This experiment is conducted on THUNDERSET. The top section compares models supervised with mouth-mesh-to-speech (mouth2s), full-mesh-to-speech (face2s) and flame-to-speech (exp2s). While it is already present in the main paper text, we include it here for completeness. The mid and bottom sections analyze the effect of weight $w_{m2s}$, with either frozen (-F) or trainable (-T) Wav2Vec2. THUNDER-T models exhibit superior lip-sync performance but it comes at the expense of reduced diversity. Furthermore, for both THUNDER-F and THUNDER-T models, increasing $w_{m2s}$, while beneficial for lip animation quality, comes at an increasing expense of of diversity. Please refer to the Sup. Video for qualitative comparison.
@@ 7.2.6 Qualitative Results
Figure 16: TFHP models trained without head pose. We
@@ 8 Architecture of Baselines
FlameFormer. Our FlameFormer architecture is based on FaceFormer. Like FaceFormer, FlameFormer is a transformer-based deterministic speech-driven animation network that takes addition categorical style condition on the input. Instead of predicting the full vertex space like the original FaceFormer does, FlameFormer predicts FLAME 3DMM coefficients $\boldsymbol{x}$. Additionally, instead of of using the autoregressive loop with a transformer decoder, FlameFormer’s decoder is a transformer encoder, which eliminates the need for the autoregressive formulation. This does not degrade performance and makes the model more efficient, as was already shown by Daněček et al. [18], which have proposed the FlameFormer baseline in their paper. Furthermore, we extend FlameFormer’s conditioning inputs - the subject identity (one-hot vector) is complemented with a one-hot vector for 8 basic emotions and three intensities.
@@ 9.1 Choice of methodology
Acquiring high quality 4D scans of sufficient scale and richness is costly, time-consuming and requires specialized hardware. Hence, recent speech-driven animation works [74, 18, 98, 95] have turned to pseudo-GT recovered from videos. This choice comes with an important decision - which face reconstruction methodolgy to use. There are several options but two of the most viable options are:
@@ 9.2 Discussion of existing FLAME regressors
SPECTRE [29] applies a paradigm similar to EMOCA but leverages a lip-reading consistency loss instead. It produces good lip-readable animations, but unfortunately often exaggarates and over-articulates, is not capable of reconstructing rich emotional expressions and the disentanglement between expression and identity is rather poor, which often makes the predicted expression vector compensate the inacuraccies in identity predictions.

### v4

@@ 1 Introduction
In summary, our contributions are: (1) A novel mesh-to-speech model that regresses sound from 3D facial animations. (2) A new form of cross-modal audio-based self-supervision via comparison of the representations of the input speech and the animation-produced output speech. (3) THUNDER, a 3D speech-driven animation method based on diffusion, capable of generating a variety of facial expressions while maintaining accurate lip-sync. The code, data annotations and models will be made publicly available at https://thunder.is.tue.mpg.de/.
@@ 2.2 Silent-video-to-speech (SVTS)
SVTS [20] was the first method to produce intelligible audio on large-scale in-the-wild datasets such as LRS2 and LRS3 [1] by regressing spectrograms from mouth crops. The spectrograms are converted to the final waveform with a pretrained vocoder[12]. Follow-up methods improve the prediction quality [48] by adding a surrogate ASR loss or predicting features from an ASR model [12] The most recent architectures are based on diffusion, producing audio often indistinguishable from real speech [11, 96]. These, however, are not suitable for a self-supervised reconstruction loop, due to the iterative nature of diffusion models. Hence, thanks to its simple, yet efficient feed-forward design and a relatively straightforward architecture and training, we base our mesh-to-speech model on Choi et al.[12].
@@ 5.3.1 Quantitative evaluation
Does the input space for M2S matter? While all models improve lip-sync metrics, mouth2speech has the strongest effect on lip-sync metrics, likely thanks to its localized effect on the mouth region (Tab. 2, experiment 1).
@@ 6 Conclusion
Acknowledgments. We thank Raja Bala, Hiro Takeda, Brandon Smith, Ganesh Iyer, Chinghang Chen, Alex Vorobiov and Betty Mohler for helpful discussions. This project was supported by an Amazon Research Award.
Disclosure: While MJB is a co-founder and Chief Scientist at Meshcapade, his contribution was performed at, and funded by, the MPG.
@@ 7.2.1 Experiments on TFHP
Consistently to the experiments on THUNDERSET, we find that THUNDER models benefit from the M2S loss, which can be observed in improvement on all lip-sync metrics. The improvement is more dramatic in THUNDER-F models, but manifests itself also in the models with trainable audio encoders (THUNDER-F). Similarly to our experiments on THUNDERSET, the application of M2S comes with a small reduction in upper-face diversity. Finally, THUNDER-T outperforms DiffPoseTalk on all lip-sync metrics. The metrics can be seen in Tab. 5, experiment 1.
@@
Finally, we train THUNDER models with the DiffPoseTalk contrastive style features as conditions. The results can be found in See Tab. 5, experiment 3. We find that conditioning with the style vector results in better lip-sync performance, lower face diversity, higher beat alignment and lower pose diversity compared to measurements in See Tab. 5, experiment 2. All of these are consistent with expectations, since passing in a style specification reduces the distribution of possible outcomes. Regardless of the above, the application of M2S has the expected effect - improved lip-sync metrics and lower diversity.
@@
A recent paper by Chae et al. [8] proposes a new self-supervised speech-mesh representation, in form of a bi-modal masked autoencoder. The architecture consists of two encoders and two decoders (one for the mesh modality and one for the audio modality). The model is trained with two loss terms - the autoencoder reconstruction losses for both modalities and the InfoNCE loss which brings the two modalities into one coherent features space. This is similar to what CLIP [66] does with images and text. The authors show that their representation can be used in training other speech-driven animation methods. They use their frozen autoencoder in training, passing the input audio and output geometry in, extracting the features for both. Then, they apply the same InfoNCE loss between the two modalities and use it to co-supervise the talking head avatar system. In this experiment, we incorporate this perceptual loss into the THUNDER training. The training loss is then given as:
@@
where $w_{perc}\mathcal{L}_{perc}$ is the perceptual term. We follow the authors and set the weight $w_{perc}=0.1$, a factor of 7 orders of magnitude lower than the vertex loss, which puts the computed losses on a similar scale and hence results in optimal effect (i.e not too strong or too weak). For completeness, we also report the new metrics proposed by Chae et al. [8] - mean temporal misalignment (MTM) and Perceptual Lip Readability Score (PLRS) along with the other metrics used in this paper. MTM measures the temporal misalignment using a Derivative Dynamic Time Warping (DDTW). SLRS is the same InfoNCE-based perceptual loss used for training, but computed by a different instance of the same architecture. The authors released two instances - one to be used as the perceptual loss and one to be used for evaluation with SLRS. We utilize both of the speech-mesh autoencoder models released by the authors accordingly. Tab. 5, experiment 4 reports the results of this experiment. In this experiment we investigate the following:
@@
(b)Does training our diffusion-based speech-driven animation method with this perceptual loss improve performance? Yes, THUNDER models trained without mesh-to-speech but with the perceptual loss improve considerably on PLRS. This is expected since PLRS is computed by a model with the same architecture as the network which provides the perceptual loss function. We also observe improvement on MTM. Furthermore, we observe improvement on our lip-sync related metrics in models with the frozen backbone (-F). Models with trainable backbones (-T) do not seem to improve on LVE, CCC, PCC or DTW when the perceptual loss is applied.
(c) Can THUNDER be trained with both mesh-to-speech and the perceptual loss. Yes. Both losses can be applied together. The mesh-to-speech loss has a strong impact on LVE, CCC, PCC, DTW and MTM, while the perceptual loss strongly affects PLRS, MTM and with -F models also improves CCC and PCC. THUNDER-F with both mesh-to-speech and the perceptual loss appears to get the best of both worlds, while THUNDER-T with mesh-to-speech only is competitive with THUNDER-T with both losses. In any case, applying the mesh-to-speech is beneficial.
@@ 7.2.3 Extended Ablation and Sensitivity Analysis
Table 7: Ablation study and sensitivity analysis. Here we analyze the effect of the mesh-to-speech loss on the talking head avatar training. This experiment is conducted on THUNDERSET. The top section compares models supervised with mouth-mesh-to-speech (mouth2s), full-mesh-to-speech (face2s) and flame-to-speech (exp2s). While it is already present in the main paper text, we include it here for completeness. The mid and bottom sections analyze the effect of weight $w_{m2s}$, with either frozen (-F) or trainable (-T) Wav2Vec2. THUNDER-T models exhibit superior lip-sync performance but it comes at the expense of reduced diversity. Furthermore, for both THUNDER-F and THUNDER-T models, increasing $w_{m2s}$, while beneficial for lip animation quality, comes at an increasing expense of diversity. Please refer to the Sup. Video for qualitative comparison.
@@ 7.2.6 Qualitative Results
Figure 16: TFHP models trained without head pose.
@@ 8 Architecture of Baselines
FlameFormer. Our FlameFormer architecture is based on FaceFormer. Like FaceFormer, FlameFormer is a transformer-based deterministic speech-driven animation network that takes addition categorical style condition on the input. Instead of predicting the full vertex space like the original FaceFormer does, FlameFormer predicts FLAME 3DMM coefficients $\boldsymbol{x}$. Additionally, instead of using the autoregressive loop with a transformer decoder, FlameFormer’s decoder is a transformer encoder, which eliminates the need for the autoregressive formulation. This does not degrade performance and makes the model more efficient, as was already shown by Daněček et al. [18], which have proposed the FlameFormer baseline in their paper. Furthermore, we extend FlameFormer’s conditioning inputs - the subject identity (one-hot vector) is complemented with a one-hot vector for 8 basic emotions and three intensities.
@@ 9.1 Choice of methodology
Acquiring high quality 4D scans of sufficient scale and richness is costly, time-consuming and requires specialized hardware. Hence, recent speech-driven animation works [74, 18, 98, 95] have turned to pseudo-GT recovered from videos. This choice comes with an important decision - which face reconstruction methodology to use. There are several options but two of the most viable options are:
@@ 9.2 Discussion of existing FLAME regressors
SPECTRE [29] applies a paradigm similar to EMOCA but leverages a lip-reading consistency loss instead. It produces good lip-readable animations, but unfortunately often exaggerates and over-articulates, is not capable of reconstructing rich emotional expressions and the disentanglement between expression and identity is rather poor, which often makes the predicted expression vector compensate the inaccuracies in identity predictions.